OpenStudy (anonymous):

I will give medal and fans to whoever can help me with this please it's urgent

OpenStudy (anonymous):

I have a math report to do on one of my family members and it says that I need to do it too by writing a system of linear equations and solve it by graphing so can someone please help me out with my part? Here's the info V V V V Problem: If you can figure out the exact coins that I found, you can have the money. If not, I'll keep it myself. Here are your clues. There are twenty coins altogether. There are only pennies and nickels. And the total value of the coins is forty-eight cents. Report: Pose the coin problem to someone else in your household, like a parent, grandparent, or older sibling…someone you think can get it! Ask them to solve the problem using whatever strategy they wish. Take notes as this family member solves the problem. Jot down their solution as well as the method they used to arrive at it. Solve the problem on your own using a different method. Think about the ways you have solved similar problems in this module. For example, your family member may have solved the problem by modeling with actual coins and using the guess-and-check strategy. You could write a system of linear equations and solve by graphing. Check your answer! Report on this process following the guidelines below. VVVV For this lesson, you must submit your work in the form of a report. The goal of the report is to provide the answer to the coin challenge and explain how you and your family member used different methods to solve it. Include general ideas as well as supporting details. Your report should be 2-4 paragraphs typed in a word processing document.

OpenStudy (anonymous):

@acxbox22 please?

OpenStudy (acxbox22):

well i mean you need a family member i guess.. what is your solution?

OpenStudy (anonymous):

@acxbox22 I will take care of the family member part I just need help with coming up with a system of linear equations and a graph to solve the equations.

OpenStudy (acxbox22):

There are twenty coins altogether. There are only pennies and nickels. And the total value of the coins is forty-eight cents. we know that information to come up with a system of equations we need two equations

OpenStudy (acxbox22):

so let x=pennies y=nickels we know together that x+y=20 since there are 20 coins all together

OpenStudy (acxbox22):

now we need another equation we know a penny is worth 1 cent and a nickel is worth 5 cents together they are worth 48 cents

OpenStudy (acxbox22):

so 1x+5y=48

OpenStudy (anonymous):

Yea we need to have one that gives how many pennies and how many nickels there is and one that solves for the total amount which is 48 cents.

OpenStudy (anonymous):

Now we have that

OpenStudy (acxbox22):

yes so now we have our two equations x+y=20 x+5y=48 do you know how to solve this system?

OpenStudy (anonymous):

If x equals the amount of pennies and y equals the amount of nickels, then lets say that we have 7 nickels. So 7 times 5 equals 35. x+y=20 x+7y=20 x+35=20 x=13

OpenStudy (anonymous):

Did I solve it right @acxbox22?

OpenStudy (acxbox22):

kind of... well use elimination